When –16 is divided by _________ the quotient is 4.

Grade:12

2 Answers

Harshit Singh
3 years ago
Dear Student

-4:
Let -16 be divided by x and the quotient is 4
So,
-16/x = 4
x = -4


Thanks
Shraddha Singh
2 years ago
Dear Student,
 
Let -16 be divided by x to give 4
 
-16÷ X = 4
-16 × 1/ X = 4
-16 = 4X
X = -16/4 = -4
 
Hope it helps!
Thank you